# FoundationGrasp: Generalizable Task-Oriented Grasping with Foundation Models

### v1

@@ Abstract
Task-oriented grasping (TOG), which refers to the problem of synthesizing grasps on an object that are configurationally compatible with the downstream manipulation task, is the first milestone towards tool manipulation. Analogous to the activation of two brain regions responsible for semantic and geometric reasoning during cognitive processes, modeling the complex relationship between objects, tasks, and grasps requires rich prior knowledge about objects and tasks. Existing methods typically limit the prior knowledge to a closed-set scope and cannot support the generalization to novel objects and tasks out of the training set. To address such a limitation, we propose FoundationGrasp, a foundation model-based TOG framework that leverages the open-ended knowledge from foundation models to learn generalizable TOG skills. Comprehensive experiments are conducted on the contributed Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset, demonstrating the superiority of FoudationGrasp over existing methods when generalizing to novel object instances, object classes, and tasks out of the training set. Furthermore, the effectiveness of FoudationGrasp is validated in real-robot grasping and manipulation experiments on a 7 DoF robotic arm. Our code, data, appendix, and video are publicly available at https://sites.google.com/view/foundationgrasp.
@@
This research is motivated by the problem of generalizable task-oriented grasping skill learning. Solving such a problem would significantly improve the robot’s level of automation and intelligence in tool manipulation for household and industrial tasks. Existing methods cannot handle unseen objects and tasks in open-world environments. To address such a problem, we propose to leverage the open-ended knowledge from foundation models to improve the generalization capability of existing TOG methods. This way, the robot can perform TOG w.r.t. unseen objects and tasks, facilitating downstream tool manipulation. The proposed method currently follows a sampling-and-evaluation pipeline. Such a design makes it highly dependent on the quality of the off-the-shelf grasp sampler. Future work aims to construct a one-stage TOG pipeline with improved robustness and efficiency. Generally, this research can be applied to diverse household and industrial scenarios involving tool manipulation, such as kitchenware cleaning and part assembly.
@@ I Introduction
Endowing intelligent robots with the ability of tool manipulation, close to the human level, opens up the potential for robots to accomplish diverse household and industrial tasks, such as part assembly and kitchenware cleaning. As an initial step towards tool manipulation, task-oriented grasping (TOG) necessitates the robot synthesizing grasps that are configurationally compatible with downstream manipulation tasks. Task-incompatible grasping will inevitably limit the success of manipulation tasks or, in some cases, pose physical risks to the user (e.g., hand over a knife by grasping the handle and exposing the sharp blade to the receiver).
@@
Despite the significance of TOG, such skills are not yet readily accessible to robots nowadays. Considering a household robot operating in environments, such as apartments and garages, it needs to handle an open set of elements, including objects (instances, classes) and tasks, most of which are novel elements unseen during training. A naive solution is to collect task-oriented grasping data for each object with respect to each task in a brute-force manner. However, such a solution is infeasible in terms of manpower and time costs in open-world environments. In practical terms, one would like to train the robot only on a limited set of objects and tasks, while the robot is able to generalize the learned TOG skills to novel elements beyond the training examples.
@@
In cognitive psychology, it is observed that two brain regions in the prefrontal cortex, which are respectively responsible for semantic and geometric reasoning, are jointly activated during cognitive processes[1, 2]. Inspired by this observation, recent works [3, 4, 5, 6, 7] have proposed incorporating semantic and geometric knowledge into TOG pipelines to improve the robots’ adaptability to different situations. For example, Murali et al.[5] contribute the largest and most diverse TOG dataset, TaskGrasp, and build a knowledge graph (KG) capturing semantic relationships between objects and tasks within the TaskGrasp dataset. In parallel, a series of works [6, 7] consider both semantic and geometric properties of pre-defined sets of objects and tasks. While the aforementioned methods have exhibited a degree of success, they experience significant performance degradation when handling novel elements out of the pre-defined sets. This can be naturally attributed to the fact that the prior knowledge is limited to a closed-set scope. While such a limitation may be acceptable under the closed-world assumption, a household robot often operates in open-world environments, thus resulting in unsatisfying TOG performance in real-world applications.
@@
The above discussions have led us to the question: How can we endow robots with open-ended knowledge about objects and tasks to learn generalizable TOG skills? Recent advancements in foundation models (e.g., Large Language Model (LLM) and Vision-Language Model (VLM)) have enabled robots to seamlessly extract and harness the open-ended knowledge embedded in these models for physically grounded tasks [8, 9, 10, 11, 12]. Motivated by this trend, we propose FoundationGrasp, a foundation model-based TOG framework. As is depicted in Figure LABEL:fig:concept, in contrast to previous works that restrict the prior knowledge to a closed-set scope, FoundationGrasp leverages the open-ended knowledge from foundation models to learn generalizable TOG skills.
@@
The pipeline of the proposed FoundationGrasp is as follows. When presented with an object (e.g., mug), a task (e.g., pour), and a language instruction (e.g., pour water from the mug), where the object and the task can be novel elements out of the training set, FoundationGrasp first prompts an LLM to generate semantic and geometric descriptions of the object and the task. Subsequently, an LLM-based semantic knowledge encoder and a VLM-based geometric knowledge encoder transform them with multi-modal sensory inputs into their feature representations. In the final stage, we draw inspiration from cognitive psychology and introduce a Transformer-based[13] task-oriented grasp evaluator with semantic and geometric branches. Evaluation on the contributed Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset demonstrates that FoundationGrasp outperforms existing TOG methods when generalizing to novel object instances, object classes, and tasks out of the training set. To validate the effectiveness in real-world applications, we further deploy FoundationGrasp on a Kinova Gen3 7 DoF robotic arm for task-oriented grasping and manipulation.
@@
This study extends a preliminary version GraspGPT [14], which showcases the potential of improving the generalization capability of TOG with an LLM in a proof-of-concept manner. Specifically, GraspGPT focuses on the semantic aspect of TOG and leverages the transferability of language to generalize to novel object classes and tasks. Building on top of GraspGPT, this extended version follows the same idea and introduces an enhanced framework, FoundationGrasp, which leverages both semantic and geometric knowledge from foundation models to learn generalizable TOG skills. More importantly, we include additional baselines, experiments, and analysis to thoroughly demonstrate the value of combining foundation models in solving TOG problems and provide insights into the limitations. An extended discussion of related work on task-agnostic grasping, task-oriented grasping, and the applications of foundation models in robotics is also presented.
@@
  • To support the training and evaluation of FoundationGrasp, we contribute the Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset.
  • A physical system is built to enable the robot to perform task-oriented grasping with novel objects and tasks guided by users’ natural language instructions.
@@ II-A Task-Agnostic Grasping
Robotic grasping has been a fundamental problem in robotics. Earlier studies optimize grasp quality to output analytic solutions satisfying stability constraints, such as force closure and form closure [15]. However, a significant drawback of the analytic approach is the need for complete object and hand models [16], hindering the generalization to novel objects.
With the rise of machine learning techniques in recent years, the field of grasp synthesis and planning has been revolutionized by data-driven methods [17, 18, 19, 20, 21] trained on various types of sensory data, including RGB-D images, depth images, and point clouds. For example, Dex-Net 2.0[17] and DeepGrasp[18] employ convolutional neural networks (CNN) to process RGB-D images and predict SE(2) planar grasp poses. PS-CNN [20] takes advantage of both analytic and data-driven approaches by performing primitive shape decomposition on depth images to generate 6 DoF grasp poses in SE(3) space. [19, 21] approach the problem of grasp synthesis on point clouds with sampling-and-evaluation pipelines. While a substantial portion of research in robotic grasping focuses on task-agnostic grasping (TAG), the goal of grasping is often beyond this realm. Ignoring the task-oriented effect on object grasping limits the application of TAG methods in many cases. FoundationGrasp complements TAG methods by predicting grasps satisfying both stability and task-compatibility constraints.
@@ II-B Task-Oriented Grasping
In terms of TOG, where the grasping problem is contextualized in manipulation tasks, the objective is to grasp the object in a manner that is configurationally compatible with the downstream manipulation task. Classical TOG methods evaluate the task-oriented grasp quality with task wrench space analysis [22]. However, they share similar limitations with earlier studies in TAG due to the inherent diversity of objects and tasks in open-world environments. Recently, data-driven methods have achieved success in solving TOG problems to some extent. Dang et al.[23] build a task-oriented grasp planner that considers partial object geometry, tactile contacts, and hand kinematic data. Yang et al.[24] expand previous works to enable TOG in object-stacking scenes. To alleviate the burden of data collection, Fang et al.[25] and Qin et al. [26] train a dedicated TOG model for each task with simulated self-supervision. Nevertheless, the aforementioned methods acquire object-task-grasp relationships solely from training data without external knowledge sources and consequently exhibit unsatisfying generalization performance.
@@
More recent works have proposed incorporating external prior knowledge into TOG pipelines. Montesano et al.[3] and Song et al. [4][27] construct semantic knowledge bases (KBs) with pre-defined sets of objects, tasks, actions, and constraints. Similarly, Ardón et al. [28] use Markov Logic Networks (MLN) to build a semantic KB, encoding objects, tasks, visual attributes, and indoor locations of 30 commonly found household objects. To enrich a KB with a broader range of objects and tasks, Murali et al.[5] contribute the largest and most diverse TOG dataset, TaskGrasp, and a resulting KG. More importantly, they propose the state-of-the-art graph-based TOG method, GCNGrasp, to generalize to objects and tasks within the KG. In addition to only incorporating semantic knowledge, a series of works [6, 7] consider both semantic and geometric properties of pre-defined sets of objects and tasks. For example, Detry et al.[6] approach the TOG problem with a geometric grasp model and a CNN-based semantic model. Similarly, in [7], a visual module that estimates object geometric properties is integrated with a probabilistic logic module. Another line of research utilizes affordance knowledge to predict task-oriented grasps. Xu et al. [29] combine affordance segmentation and keypoint detection to guide task-oriented grasp prediction. Liu et al.[30] propose a novel representation of task-oriented grasp contexts with part affordances, part material, object states, and tasks.
@@
While these methods have demonstrated the generalization capability to some extent, they experience significant performance degradation when handling novel objects and tasks out of the pre-defined sets. This phenomenon can be explained by the fact that these methods limit the prior knowledge to a closed-set scope. To learn generalizable TOG skills, we propose to leverage open-ended knowledge from foundation models in this work.
@@ II-C Foundation Models in Robotics
Foundation models, such as LLMs and VLMs, encode open-ended knowledge about the world and have exhibited impressive capabilities on a wide range of visual and linguistic tasks. In robotics, such knowledge is beneficial for physically grounded tasks, such as tabletop manipulation[8, 9], navigation[10, 11], and mobile manipulation[12]. While most of these methods primarily rely on foundation models for task planning and scene understanding, FoundationGrasp grounds the open-ended knowledge to the grasp synthesis process.
Regarding robotic grasping, LERF-TOGO[31] builds a Language Embedded Radiance Fields (LERF) with features distilled from a VLM to produce a 3D relevancy heatmap given an object part name. LAN-Grasp [32] first prompts an LLM to output an object part name appropriate for grasping and uses a VLM to identify the corresponding part in the image. These two methods focus on solving “where to grasp” and ignore “how to grasp”. FoundationGrasp optimizes 6 DoF task-oriented grasp poses to jointly address “where to grasp” and “how to grasp”. GraspCLIP[33] and ATLA[34] extract knowledge from foundation models for manipulation and grasping. However, they are limited to 2D planar grasping. In contrast, FoudationGrasp studies TOG in 3D space.
III Problem Formulation
$g^{*}=\pi(\gamma)=\pi(X_{o},I_{o},L)$ (1)
where $X_{o}\in\mathbb{R}^{N\times 3}$ and $I_{o}\in\mathbb{R}^{H\times W\times 3}$ are the (partial) point cloud and the RGB image of the object instance $o$, respectively. $N$ is the number of points. $H$ and $W$ are the height and width of the image. $L$ is a natural language instruction specifying an object class $c$ and a task $t$, such as “pour water with the mug”. $g^{*}$ is a 6 DoF grasp pose of a parallel-jaw gripper represented by $(R,T)$, where $R\in SO(3)$ represents the 3D orientation and $T\in\mathbb{R}^{3}$ represents the 3D translation.
Figure 1: An overview of FoundationGrasp framework: The pipeline consists of (1) data generation, (2) multi-modal feature representation, and (3) task-oriented grasp evaluation. When presented with a language instruction, FoundationGrasp first prompts an LLM to generate semantic and geometric descriptions of the object and the task. A web-based image retrieval module crowdsources images from the Internet. Subsequently, an LLM-based semantic knowledge encoder and a VLM-based geometric knowledge encoder transform them with multi-modal sensory inputs into their feature representations. In the final stage, a Transformer-based task-oriented grasp evaluator with semantic and geometric branches evaluates the task compatibility of each grasp candidate.
IV FoundationGrasp
IV-A Overview
An overview of the proposed FoundationGrasp framework is depicted in Figure 1. The pipeline consists of three stages: (1) data generation, (2) multi-modal feature representation, and (3) task-oriented grasp evaluation. In Section IV-B, FoundationGrasp generates linguistic data (i.e., language descriptions and instructions), visual data (i.e., images), and grasp data (i.e., grasp poses) based on $o$, $c$, and $t$. The second stage described in Section IV-C transforms multi-modal data into their latent space feature representations. Finally, in Section IV-D, a Transformer-based task-oriented grasp evaluator computes the task-compatibility score $S$ of each candidate grasp pose $g$.
IV-B Data Generation
The data generation stage constructs the LaViA-TaskGrasp dataset by augmenting the TaskGrasp dataset with linguistic and visual data. The original TaskGrasp dataset consists of $K_{o}$ object instances, spanning $K_{c}$ object classes and $K_{t}$ tasks. The data generation pipeline consists of four phases: (1) language description generation, (2) language instruction generation, (3) web-based image retrieval, and (4) grasp candidate generation.

### v2

@@ Abstract
Task-oriented grasping (TOG), which refers to synthesizing grasps on an object that are configurationally compatible with the downstream manipulation task, is the first milestone towards tool manipulation. Analogous to the activation of two brain regions responsible for semantic and geometric reasoning during cognitive processes, modeling the intricate relationship between objects, tasks, and grasps necessitates rich semantic and geometric prior knowledge about these elements. Existing methods typically restrict the prior knowledge to a closed-set scope, limiting their generalization to novel objects and tasks out of the training set. To address such a limitation, we propose FoundationGrasp, a foundation model-based TOG framework that leverages the open-ended knowledge from foundation models to learn generalizable TOG skills. Extensive experiments are conducted on the contributed Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset, demonstrating the superiority of FoundationGrasp over existing methods when generalizing to novel object instances, object classes, and tasks out of the training set. Furthermore, the effectiveness of FoundationGrasp is validated in real-robot grasping and manipulation experiments on a 7-DoF robotic arm. Our code, data, appendix, and video are publicly available at https://sites.google.com/view/foundationgrasp.
@@
This research is motivated by the challenge of generalizable task-oriented grasping skill learning. Solving such a challenge could significantly improve the robot’s level of automation and intelligence in tool manipulation for household and industrial tasks. Existing methods struggle with handling unseen objects and tasks in dynamic, open-world environments. To overcome this limitation, we propose to leverage the open-ended knowledge from foundation models to improve the generalization capabilities of existing TOG methods. This way, the robot can perform TOG w.r.t. unseen objects and tasks, facilitating downstream tool manipulation. Overall, this research has broad applicability to various scenarios involving tool manipulation, such as cleaning kitchenware and assembling parts in industrial contexts.
@@ I Introduction
Endowing intelligent robots with tool manipulation capabilities, approaching human levels, could significantly expand their potential to perform a variety of household and industrial tasks, such as part assembly and kitchenware cleaning. As a fundamental step towards tool manipulation, task-oriented grasping (TOG) necessitates robots synthesizing grasps configurationally compatible with the downstream manipulation task. Task-incompatible grasping will inevitably limit the success of manipulation tasks or, in some cases, pose physical risks to the user (e.g., handover a knife by grasping the handle and exposing the sharp blade to the receiver).
@@
Despite the significance of TOG, such skills are not yet available to robots. Considering a household robot operating in open-world environments, such as apartments and garages, it needs to handle an open set of elements, including novel objects (instances, classes) and tasks unseen during training. A straightforward approach is to collect TOG data for each object with respect to each task in a brute-force manner. However, this approach is impractical due to the immense manpower and time required. In practical terms, one would expect training the robot only on a limited set of objects and tasks, while the robot is capable of generalizing the learned TOG skills to novel elements beyond the training examples.
@@
Prior work in TOG has approached this expectation through two main paths: the analytic approach and the data-driven approach. The former evaluates the task-oriented grasp quality with task wrench space analysis [1]. However, it assumes the availability of full object and hand models, which makes it less suitable for handling novel elements in open-world environments. Consequently, recent research has shifted towards the data-driven approach [2, 3, 4, 5], which holds promise for better generalization capability. Especially, inspired by the observation that two brain regions in the prefrontal cortex, responsible for semantic and geometric reasoning, are jointly activated during cognitive processes[6, 7], more recent data-driven TOG methods [8, 9, 10, 11, 12] have integrated semantic and geometric knowledge into TOG pipelines to enhance adaptability. For example, Murali et al.[10] contribute the largest and most diverse TOG dataset, TaskGrasp, and build a knowledge graph (KG) capturing semantic relationships between objects and tasks within the TaskGrasp dataset. Concurrently, several studies [11, 12] consider both semantic and geometric properties of pre-defined sets of objects and tasks. Despite the successes of these methods, they struggle to generalize to novel elements out of the pre-defined sets. This limitation can be naturally attributed to the closed-set nature of the prior knowledge used. While such a limitation may be acceptable under the closed-world assumption, it results in sub-optimal TOG performance for household robots operating in open-world environments.
@@
The preceding discussion raises a critical question: How can we endow robots with open-ended knowledge about objects and tasks to learn generalizable TOG skills? Recent advancements in foundation models, such as Large Language Models (LLMs) and Vision-Language Models (VLMs), have enabled robots to seamlessly extract and harness the open-ended knowledge embedded in these models for physically grounded tasks [13, 14, 15, 16, 17]. Motivated by this trend, we propose FoundationGrasp, a foundation model-based TOG framework. As depicted in Figure LABEL:fig:concept, in contrast to previous works that restrict the prior knowledge to a closed-set scope, FoundationGrasp leverages the open-ended knowledge from foundation models to learn generalizable TOG skills.
@@
The pipeline of the proposed FoundationGrasp operates as follows. When presented with an object (e.g., mug), a task (e.g., pour), and a language instruction (e.g., pour water from the mug), where the object and the task can be novel elements outside the training set, FoundationGrasp begins by prompting an LLM to generate semantic and geometric descriptions of the object and the task. Subsequently, an LLM-based semantic knowledge encoder and a VLM-based geometric knowledge encoder transform them with multi-modal sensory inputs into their feature representations. In the final stage, we draw inspiration from cognitive psychology and introduce a Transformer-based[18] task-oriented grasp evaluator with semantic and geometric branches. Evaluation on the contributed Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset demonstrates that FoundationGrasp outperforms existing TOG methods when generalizing to novel object instances, object classes, and tasks beyond the training set. To validate the effectiveness in real-world applications, we further deploy FoundationGrasp on a Kinova Gen3 7-DoF robotic arm for task-oriented grasping and manipulation.
@@
This study builds on the preliminary work presented in GraspGPT [19], which showcases the potential of improving the generalization capability of TOG with an LLM in a proof-of-concept manner. GraspGPT primarily addresses the semantic aspect of TOG by leveraging language transferability to establish semantic connections between novel and familiar linguistic concepts (i.e., object class labels and task labels). However, semantic understanding alone is insufficient. Geometric understanding is also crucial for TOG since it enables the robot to visually identify different object parts and local geometries suitable for various tasks and affordances. For example, “spoon” and “saucepan” are semantically different but geometrically similar. Solely using semantic knowledge struggles to relate these concepts. However, both objects share the same geometric decomposition: a long handle plus a functional head. Building on this intuition, the extended version, FoundationGrasp, leverages both semantic and geometric knowledge from foundation models to learn generalizable TOG skills. Moreover, we include additional baselines, experiments, and studies to thoroughly demonstrate the value of incorporating foundation models in solving TOG problems and provide insights into the limitations. An extended discussion of related work on task-agnostic grasping, task-oriented grasping, and the applications of foundation models in robotics is also presented.
@@
  • Built on top of the public TOG benchmark TaskGrasp dataset, we contribute a multi-modal TOG dataset, Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset, to support the training and evaluation of generalizable TOG methods.
@@ II-A Task-Agnostic Grasping
Grasping has been a fundamental problem in robotics. Earlier studies optimize grasp quality to output analytic solutions satisfying stability constraints, such as force closure and form closure [20]. However, a significant drawback of the analytic approach is its reliance on complete object and hand models [21], which impedes its ability to generalize to novel objects.
With the rise of machine learning techniques in recent years, the field of grasp synthesis and planning has been revolutionized by data-driven methods [22, 23, 24, 25, 26] trained on various types of sensory data, including RGB-D images, depth images, and point clouds. For example, Dex-Net 2.0[22] and DeepGrasp[23] employ convolutional neural networks (CNNs) to process RGB-D images and predict SE(2) planar grasp poses. PS-CNN [25] takes advantage of both analytic and data-driven approaches by performing primitive shape decomposition on depth images to generate 6-DoF grasp poses in SE(3) space. [24, 26] approach the problem of grasp synthesis on point clouds with sampling-and-evaluation pipelines. While a substantial portion of research in robotic grasping focuses on task-agnostic grasping (TAG), the goal of grasping is often beyond this realm. Ignoring the task-oriented effect on object grasping limits the application of TAG methods in many cases. FoundationGrasp complements TAG methods by predicting grasps satisfying both stability and task-compatibility constraints.
@@ II-B Task-Oriented Grasping
In terms of TOG, where the grasping problem is contextualized in manipulation tasks, the objective is to grasp the object in a manner that is configurationally compatible with the downstream manipulation task. Classical TOG methods evaluate the task-oriented grasp quality with task wrench space analysis [1]. However, they share similar limitations with earlier studies in TAG due to the inherent diversity of objects and tasks in open-world environments. Recently, data-driven methods have achieved success in solving TOG problems to some extent. Dang et al.[2] build a task-oriented grasp planner that considers partial object geometry, tactile contacts, and hand kinematic data. Yang et al.[3] expand previous works to enable TOG in object-stacking scenes. To alleviate the burden of data collection, Fang et al.[4] and Qin et al. [5] train task-dependent TOG models with simulated self-supervision. Nevertheless, the aforementioned methods derive object-task-grasp relationships solely from training data without external knowledge sources and consequently exhibit limited generalization performance. Diffusion models have emerged as potential tools for TOG [27, 28, 29]. However, their generalization capabilities remain questionable.
@@
More recent works have incorporated external prior knowledge into TOG pipelines. Montesano et al.[8] and Song et al. [9][30] construct semantic knowledge bases (KBs) with pre-defined sets of objects, tasks, actions, and constraints. Similarly, Ardón et al. [31] use Markov Logic Networks (MLN) to build a semantic KB, encoding objects, tasks, visual attributes, and indoor locations of 30 commonly found household objects. To enrich a KB with a broader range of objects and tasks, Murali et al.[10] contribute the largest and most diverse TOG dataset, TaskGrasp, and a resulting KG. More importantly, they propose the state-of-the-art graph-based TOG method, GCNGrasp, to generalize to objects and tasks within the KG. In addition to only incorporating semantic knowledge, a series of works [11, 12] consider both semantic and geometric properties of pre-defined sets of objects and tasks. For example, Detry et al.[11] approach the TOG problem with a geometric grasp model and a CNN-based semantic model. Similarly, in [12], a visual module that estimates object geometric properties is integrated with a probabilistic logic module. Another line of research utilizes affordance knowledge to predict task-oriented grasps[32, 33, 34]. Xu et al. [32] combine affordance segmentation and keypoint detection to guide task-oriented grasp prediction. Liu et al.[34] propose a novel TOG representation with part affordances, part material, object states, and tasks.
@@
While these methods have demonstrated the generalization capability to some extent, they suffer from generalization to novel objects and tasks beyond pre-defined sets. This phenomenon can be explained by the fact that these methods limit the prior knowledge to a closed-set scope. To learn generalizable TOG skills, we propose to leverage open-ended knowledge from foundation models in this work.
@@ II-C Foundation Models in Robotics
Foundation models, such as LLMs and VLMs, encode open-ended knowledge about the world and have exhibited impressive capabilities on a wide range of visual and linguistic tasks. In robotics, such knowledge is beneficial for physically grounded tasks, such as tabletop manipulation[13, 14], navigation[15, 16], and mobile manipulation[17]. While most of them primarily rely on foundation models for task planning and scene understanding, FoundationGrasp grounds the open-ended knowledge to the grasp synthesis process.
Regarding robotic grasping, LERF-TOGO[35] builds a Language Embedded Radiance Fields (LERF) with features distilled from a VLM to produce a 3D relevancy heatmap given an object part name. LAN-Grasp [36] first prompts an LLM to output an object part name appropriate for grasping and uses a VLM to identify the corresponding part in the image. These methods[35, 36, 37] focus on solving “where to grasp” and ignore “how to grasp”. FoundationGrasp optimizes 6-DoF task-oriented grasp poses to jointly address “where to grasp” and “how to grasp”. GraspCLIP[38] and ATLA[39] extract knowledge from foundation models for manipulation and grasping. However, they are limited to 2D planar grasping. In contrast, FoundationGrasp studies TOG in 3D space.
@@
III Problem Formulation
$g^{*}=\pi(\gamma)=\pi(X_{o},I_{o},L)$ (1)
where $X_{o}\in\mathbb{R}^{N\times 3}$ and $I_{o}\in\mathbb{R}^{H\times W\times 3}$ are the (partial) point cloud and the RGB image of the object instance $o$, respectively. $N$ is the number of points. $H$ and $W$ are the height and width of the image. $L$ is a natural language instruction specifying an object class $c$ and a task $t$, such as “pour water with the mug”. $g^{*}$ is a 6-DoF grasp pose of a parallel-jaw gripper represented by $(R,T)$, where $R\in SO(3)$ represents the 3D orientation and $T\in\mathbb{R}^{3}$ represents the 3D translation.
Figure 1: An overview of FoundationGrasp framework: The pipeline consists of (1) knowledge generation, (2) multi-modal feature representation, and (3) task-oriented grasp evaluation. When presented with a language instruction, FoundationGrasp first prompts an LLM to generate semantic and geometric descriptions of the object and the task. A web-based image retrieval module crowdsources images from the Internet. Subsequently, an LLM-based semantic knowledge encoder and a VLM-based geometric knowledge encoder transform them with multi-modal sensory inputs into their latent space feature representations. In the final stage, a Transformer-based task-oriented grasp evaluator with semantic and geometric branches evaluates the task compatibility of each grasp candidate.
IV FoundationGrasp
IV-A Overview
An overview of the proposed FoundationGrasp framework is depicted in Figure 1. The pipeline consists of three stages: (1) knowledge generation, (2) multi-modal feature representation, and (3) task-oriented grasp evaluation. In Section IV-B, FoundationGrasp generates linguistic data, visual data, and grasp data based on $o$, $c$, and $t$ with foundation models or pre-trained models. The second stage described in Section IV-C transforms multi-modal data into their latent space feature representations. Finally, in Section IV-D, a Transformer-based task-oriented grasp evaluator computes the task-compatibility score $S$ of each candidate grasp pose $g$.
IV-B Knowledge Generation
The knowledge generation stage constructs the LaViA-TaskGrasp dataset by augmenting the TaskGrasp dataset with linguistic and visual data. The original TaskGrasp dataset consists of $K_{o}$ object instances, spanning $K_{c}$ object classes and $K_{t}$ tasks. The knowledge generation pipeline consists of four phases: (1) language description generation, (2) language instruction generation, (3) web-based image retrieval, and (4) grasp candidate generation.